# Designing AI-Agents with Personalities: A Psychometric Approach

### v3

@@ 1 Introduction
The emergence of large language models (LLMs) has revolutionized our approach to simulating human-like behaviors and communication. LLMs are increasingly deployed across diverse research fields to mimic human behaviors [120, 121]. In psychology, LLMs are used to study cognitive processes, measure personality, and understand psychological constructs [84, 86, 104, 73]. Sociologists employ these models to explore social bias and behavior [93, 124, 101], while economists and political scientists utilize them to analyze economic processes and political leanings [71, 85]. These broad applications underscore the growing interconnection between artificial intelligence and the social and behavioral sciences.
Traditional human subjects research methods, while standard in social and behavioral studies, face significant challenges including ethical constraints, logistical hurdles, and resource limitations [107, 79]. Recent advancements in LLMs offer valuable complementary tools to address these challenges [64, 79]. Researchers can use LLMs to create AI-Agents, which can mimic responses from human participants, reducing logistical and financial burdens [117]. These AI-Agents operate continuously, simulate diverse demographic responses, and can be deployed in scenarios that might pose ethical risks to human participants [67, 65, 69].
The integration of AI-Agents in social and behavioral science research serves multiple practical purposes, complementing rather than replacing human participants. They can be used for pilot testing studies, allowing researchers to refine experimental designs and identify potential issues before investing in full-scale human participant studies. Additionally, AI-Agents can independently replicate findings from human-subjects data, enhancing the robustness and generalizability of research outcomes [90, 124]. This approach provides a complementary method for validation and exploration, contributing to the overall rigor and efficiency of social and behavioral science research. By offering a novel approach to corroborate and extend findings obtained through traditional human participant studies, AI-Agents represent an advancement in social science research methodology, enhancing inclusivity and efficiency while preserving the essential depth of human-centered understanding critical in the field.
@@ 1.1 Problems with Previous Approaches for Assigning Personas to AI-Agents
Among the myriad traits that can be incorporated into AI-Agent personas, personalities stand out as among the most intuitive characteristics to simulate. Personalities encapsulate a spectrum of human behaviors and tendencies crucial for the prediction of life outcomes ranging from academic and career to health and socioeconomic outcomes [111, 113, 110] to a multitude of interaction-based phenomena [82, 78]. Despite this potential and the capacity of AI-Agents to advance personality research, existing methodologies for implementing personality in AI-Agents remain limited, typically falling into three distinct approaches.
@@
First, personas are assigned with simple personality adjectives in prompts. For example, [86] assigned personalities to an AI-Agent by telling it “You are a character who is introverted, antagonistic, conscientious, emotionally stable, and open to experience” This type of approach holds a binary view, with either high-low or presence-absence of a trait, which contradicts established empirical evidence that personality traits exist on continuous spectra [68, 94, 126].
@@
Second, personas may be assigned through demographic descriptions and personal preferences [101, 109, 70]. For instance, [101] used narrative descriptions such as “John Lin is a pharmacy shopkeeper at the Willow Market and Pharmacy who loves to help people.” This approach relies heavily on stereotypical information extrapolated from the LLM’s training corpora, where names and occupations may trigger implicit assumptions (e.g., associating "John Lin" with Asian ethnicity and pharmaceutical expertise). This approach, however, poses difficulties in evaluating the full scope of stereotypical implications (i.e., potential confounds) and fails to provide the granularity and precision necessary for social and behavioral science studies, particularly when the degree of trait expression is crucial for creating AI-Agents with specific personality profiles.
@@
Third, personas may be embedded at the parameter level through methods like fine-tuning or direct weight editing. In one approach, researchers fine-tune an LLM on an individual-specific or group-specific text corpus, producing a distinct model persona for each dataset [114, 92]. Alternatively, knowledge editing techniques modify the model’s internal weights or latent representations to encode desired traits. For example, recent work identifies “persona vectors”, which are the specific directions in a model’s activation space associated with character traits [75, 88]. Applying such a vector inside the model can induce a target personality (e.g. making the model more sycophantic) without full retraining. While these parameter-level interventions alter behavior more fundamentally than in-context prompts, they still lack fine-grained control over how traits manifest. Moreover, locating and adjusting the correct parameters or latent directions for a given persona is technically challenging. The process often involves complex optimization and substantial computational resources, demanding deep expertise and making it less accessible to researchers without specialized backgrounds or institutional support.
@@
These approaches offer a superficial semblance of personality but suffer from critical limitations. Relying on stereotypes underlying traits (e.g., naive) or roles (e.g., policing officer) does not provide the precise control necessary for rigorous social and behavioral research and does not reflect the continuity and complexity inherent in human personalities [77].
@@ 1.2 Psychometric Approach for Assigning LLM-Agents Persona
In this article, we propose creating AI-Agents with personalities through a psychometric approach by leveraging Big Five Personality theory. Historically, Big Five personality theory was developed based on the Lexical Hypothesis, which states that personality characteristics fundamental to humans have become a part of human language, and the most important characteristics are encoded by a single word [74]. Based on the Lexical Hypothesis, scholars selected personality-related words from the English dictionary [66], refined the list of words [98, 99, 119], had participants rate themselves against those descriptors, applied factor analysis to reduce the dimensionality of the data, and eventually identified the Big Five Personality traits [95, 83, 81, 102]: Openness (O), Conscientiousness (C), Extraversion (E), Agreeableness (A) and Neuroticism (N).
Over the past decades, many psychometric scales [^1] The terms “scale” and “test” are used interchangeably to refer to psychometric instruments for assessing personality traits. measuring the Big Five traits have been developed, notably the Big Five Inventory [[, BFI;]]john_big_1991, Big Five Inventory-2 [[, BFI2;]]soto_next_2017, Mini-Markers [108], the Big Five Aspects Scale [[, BFAS;]]deyoung_between_2007 and the NEO Five-Factor Inventory [[, NEO-FFI;]]costa_neo_1989. Despite having different items, these psychological measurements demonstrated high reliability, convergent validity and predictive validity in samples from diverse backgrounds [97, 96, 103]. Specifically, these psychological measurements of the Big Five can significantly predict life outcomes, such as educational achievement, socioeconomic status, health, interpersonal relationships with peers and family, and more, indicating that personality traits can indeed forecast a range of individual, interpersonal and societal behaviors effectively [111, 100, 105].
@@
Bringing psychometric principles to design and evaluate has practical benefits [118, 123]. Integrating well-established psychometric tests (i.e., high reliability and validity) into the design of AI-Agents could ensure that the personality traits assigned to AI-Agents are stable and reflective of their programmed characteristics, enhancing the credibility and robustness of findings in applications where the consistency of AI-Agents performance is critical. In addition, as psychometric tests are designed to discern different levels of expression related to psychological traits, we can reverse engineer that variability into AI-Agents to create diversity, which is necessary for simulating real-world settings. Additionally, free from external confounds, this approach allows for precise and fine-grained manipulation over the psychological constructs of interest, allowing nuanced understanding and refinement of AI-Agents’ personalities for tailored applications.
@@ 2.1 Methods
Our analysis incorporated content from widely recognized Big Five tests, such as the BFI [87], BFI-2 [112], Mini-Markers [108], BFAS [80] and NEO-FFI [76]. We extracted domain-specific content (i.e., test items) from these tests and processed it through OpenAI’s advanced text-embedding model (text-embedding-3-large). Text embeddings are commonly used in natural language processing; they allow researchers to represent text data (such as words, phrases, sentences, or even entire documents) in a numeric format (i.e., vector) that quantifies the semantic relationships between words, phrases, or entire documents. In the context of our study, each personality test item is converted into a high-dimensional vector (3072 dimensions for the model we used) that captures its semantic meaning.
@@ 2.2 Embeddings Techniques
We applied t-SNE [115] to item embeddings to create a two-dimensional visualization of their similarity structure. t-SNE is a non-linear dimensionality reduction method that maps high-dimensional data to a low-dimensional space while preserving local relationships between data points. In our context, this provides an intuitive visual map of semantic neighborhoods among personality test items, complementing cosine similarity analysis. Items from the same Big Five domain tend to cluster together in the plot, and cross-domain proximities become easier to spot visually. In our t-SNE visualizations, items that cluster together can be interpreted as measuring similar constructs [^2] Because t-SNE is intended for visualization rather than formal measurement modeling, the plots are used here descriptively to illustrate patterns in the embedding space; the axes have no intrinsic psychological meaning..
@@ 3.1.1 Method
We included four prompt conditions for creating AI-Agents with personalities (see Table 1). Two conditions involved using the BFI-2 in either Likert or Expanded format, which we refer to as BFI-2-Likert and BFI-2-Expanded, respectively. The BFI-2-Likert is the original BFI-2 by [112], in which participants indicated their agreement with statements measuring the Big Five traits. The BFI-2-Likert, however, is susceptible to response bias and careless responding; therefore, [125] developed the BFI-2-Expanded version, replacing the agree-disagree response options with complete sentences. For example, for the Expanded format, the “strongly agree” option for the item “Is outgoing, sociable” is replaced with “I am very outgoing, sociable”. For prompt engineering, the Expanded format produces prompts that sound more natural and straightforward. As shown in Table 1, the BFI-2-Likert prompt requires an additional instruction for interpreting numeric numbers indicating agreement levels, whereas the BFI-2-Expanded prompt simply describes the trait levels in complete sentences.
We also included two baseline prompt conditions (Simple-Binary and Elaborated-Binary), designed to mimic the approach of previous studies that used simple adjectives to create AI-Agent personas [86, 101, 109]. In the Simple-Binary condition, prompts were binary Big Five statements indicating whether the AI-Agent was high or low on each trait. The Elaborated-Binary condition added descriptors such as “outgoing” and “compassionate” to further elaborate on the traits. Compared with these baseline conditions, we aim to show that the BFI-2-Likert or BFI-2-Expanded condition create AI-Agents with more nuanced personalities and aligns better with human responses.
@@
### Your Personality ### I am very outgoing, sociable. I am very compassionate almost always soft-hearted. I am fairly organized. I am somewhat relaxed, handle stress somewhat well. I have some artistic interests. I am quite assertive. I am very respectful almost always treat others with respect. I am often lazy. I stay very optimistic after experiencing a setback. I am curious about few things. I often feel excited or eager. I rarely find fault with others. I am very dependable, steady. I am fairly moody often have up and down mood swings. I am fairly inventive, often find clever ways to do things. I am rarely quiet. I feel a great deal of sympathy for others. I am not particularly systematic rarely keep things in order. I am sometimes tense. I am very much fascinated by art, music or literature. I am fairly submissive, often act as a follower. I rarely start arguments with others. I have a fair amount of difficulty getting started on tasks. I feel very secure, comfortable with self. I typically seek out intellectual, philosophical discussions. I am somewhat more active than other people. I have a very forgiving nature. I am often careless. I am fairly emotionally stable quite hard to upset. I have little creativity. I am often shy, introverted. I am very helpful and unselfish with others. I sometimes keep things neat and tidy. I worry quite a lot. I value art and beauty quite a bit. I find it fairly easy to influence people. I am sometimes rude to others. I am fairly efficient, get things done fairly quickly. I almost never feel sad. I am not particularly complex rarely a deep thinker. I am almost always full of energy. I am quite trusting of others’ intentions. I am fairly reliable, can usually be counted on. I usually keep my emotions under control. I have a bit of difficulty imagining things. I am very talkative. I am very warm and caring. I rarely leave a mess usually clean up. I often feel anxious or afraid. I think poetry and plays are fairly interesting. I strongly prefer to take charge. I am very polite and courteous to others. I am fairly persistent, usually work until the task is finished. I sometimes feel depressed and blue. I have little interest in abstract ideas. I show a lot of enthusiasm. I almost always assume the best about people. I often behave irresponsibly. I am not at all temperamental almost never get emotional. I am not at all original almost never come up with new ideas.
@@
To examine alignment between AI-Agents responses with those from human participants, we repurposed data collected by [112] wherein participants (N = 438) responded to multiple Big Five tests, including: (1) BFI-2: a sophisticated 60-item Likert test designed to capture the comprehensive hierarchical structure of personality, where each domain has three facets and (2) Mini-Markers: a straightforward, 40-item test consisting of phenotypic Big Five descriptive adjectives [108]. Participants’ BFI2 responses will serve as the training (input) data, while their Mini-Markers responses (output) will be used for validation. The human data [112] will serve as a reference for our comparison.
@@
Using the four types of prompts in Table 1, we created AI-Agents across five popular LLMs, including 1) GPT-3.5, 2) GPT-4, 3) GPT-4o, 4) DeepSeek-V3, and 5) Llama-3.3-70B-Instruct. For each LLMs and each prompt condition, we initiated 438 AI-Agents, each corresponding to a human participant’s data collected from [112]. More specifically, for the BFI-2-Likert and BFI-2-Expanded conditions, AI-Agents were created by inputting prompts that matched the BFI-2 scores of the corresponding human participants. For the Baseline-Simple and Baseline-Elaborated conditions, participants’ trait scores were dichotomized as high (average score >2.5>2.5) or low (average score <2.5<2.5) for each of the Big Five traits, and AI-Agents were then created using these dichotomized scores.
After creating AI-Agents for each LLM and prompt condition, we prompted them to complete the Mini-Markers test, which are 40 adjectives describing the Big Five traits [108]. The full prompt for this task is provided in the Appendix A. In this prompt, AI-Agents were instructed to consider the Mini-Marker items carefully based on their assigned personality (see Appendix).

### v4

@@ 1 Introduction
The emergence of large language models (LLMs) has revolutionized our approach to simulating human-like behaviors and communication. LLMs are increasingly deployed across diverse research fields to mimic human behaviors [xi_rise_2023, xu_ai_2024]. In psychology, LLMs are used to study cognitive processes, measure personality, and understand psychological constructs [hagendorff_human_like_2023, jiang_personallm_2024, rathje2024gpt, binz2025foundation]. Sociologists employ these models to explore social bias and behavior [lucy_gender_2021, zhang2025evolvingcollectivecognitionhumanagent, park_generative_2023], while economists and political scientists utilize them to analyze economic processes and political leanings [bang-etal-2024-measuring, hartmann2023political]. These broad applications underscore the growing interconnection between artificial intelligence and the social and behavioral sciences.
Traditional human subjects research methods, while standard in social and behavioral studies, face significant challenges including ethical constraints, logistical hurdles, and resource limitations [salganik_bit_2019, demszky_using_2023]. Recent advancements in LLMs offer valuable complementary tools to address these challenges [agnew_illusion_2024, demszky_using_2023]. Researchers can use LLMs to create AI-Agents, which can mimic responses from human participants, reducing logistical and financial burdens [wang2021want]. These AI-Agents operate continuously, simulate diverse demographic responses, and can be deployed in scenarios that might pose ethical risks to human participants [argyle_out_2023, aher_using_2023, bai2022constitutional].
The integration of AI-Agents in social and behavioral science research serves multiple practical purposes, complementing rather than replacing human participants. They can be used for pilot testing studies, allowing researchers to refine experimental designs and identify potential issues before investing in full-scale human participant studies. Additionally, AI-Agents can independently replicate findings from human-subjects data, enhancing the robustness and generalizability of research outcomes [kozlowski2024silicosociologyforecastingcovid19, zhang2025evolvingcollectivecognitionhumanagent]. This approach provides a complementary method for validation and exploration, contributing to the overall rigor and efficiency of social and behavioral science research. By offering a novel approach to corroborate and extend findings obtained through traditional human participant studies, AI-Agents represent an advancement in social science research methodology, enhancing inclusivity and efficiency while preserving the essential depth of human-centered understanding critical in the field.
@@ 1.1 Problems with Previous Approaches for Assigning Personas to AI-Agents
Among the myriad traits that can be incorporated into AI-Agent personas, personalities stand out as among the most intuitive characteristics to simulate. Personalities encapsulate a spectrum of human behaviors and tendencies crucial for the prediction of life outcomes ranging from academic and career to health and socioeconomic outcomes [soto_how_2019, stewart2022finer, soldz1999big] to a multitude of interaction-based phenomena [furnham1999personality, tapus2014towards]. Despite this potential and the capacity of AI-Agents to advance personality research, existing methodologies for implementing personality in AI-Agents remain limited, typically falling into three distinct approaches.
@@
First, personas are assigned with simple personality adjectives in prompts. For example, [jiang_personallm_2024] assigned personalities to an AI-Agent by telling it “You are a character who is introverted, antagonistic, conscientious, emotionally stable, and open to experience” This type of approach holds a binary view, with either high-low or presence-absence of a trait, which contradicts established empirical evidence that personality traits exist on continuous spectra [asendorpf2006typeness, marcus2006antisocial, zrari2024assessing].
@@
Second, personas may be assigned through demographic descriptions and personal preferences [park_generative_2023, serapio2023personality, bai2025irotehumanliketraitselicitation]. For instance, [park_generative_2023] used narrative descriptions such as “John Lin is a pharmacy shopkeeper at the Willow Market and Pharmacy who loves to help people.” This approach relies heavily on stereotypical information extrapolated from the LLM’s training corpora, where names and occupations may trigger implicit assumptions (e.g., associating "John Lin" with Asian ethnicity and pharmaceutical expertise). This approach, however, poses difficulties in evaluating the full scope of stereotypical implications (i.e., potential confounds) and fails to provide the granularity and precision necessary for social and behavioral science studies, particularly when the degree of trait expression is crucial for creating AI-Agents with specific personality profiles.
@@
Third, personas may be embedded at the parameter level through methods like fine-tuning or direct weight editing. In one approach, researchers fine-tune an LLM on an individual-specific or group-specific text corpus, producing a distinct model persona for each dataset [tan2025democratizinglargelanguagemodels, liu2024llm]. Alternatively, knowledge editing techniques modify the model’s internal weights or latent representations to encode desired traits. For example, recent work identifies “persona vectors”, which are the specific directions in a model’s activation space associated with character traits [chen2025personavectorsmonitoringcontrolling, kim2025linear]. Applying such a vector inside the model can induce a target personality (e.g. making the model more sycophantic) without full retraining. While these parameter-level interventions alter behavior more fundamentally than in-context prompts, they still lack fine-grained control over how traits manifest. Moreover, locating and adjusting the correct parameters or latent directions for a given persona is technically challenging. The process often involves complex optimization and substantial computational resources, demanding deep expertise and making it less accessible to researchers without specialized backgrounds or institutional support.
@@
These approaches offer a superficial semblance of personality but suffer from critical limitations. Relying on stereotypes underlying traits (e.g., naive) or roles (e.g., policing officer) does not provide the precise control necessary for rigorous social and behavioral research and does not reflect the continuity and complexity inherent in human personalities [cummings_introduction_2019].
@@ 1.2 Psychometric Approach for Assigning LLM-Agents Persona
In this article, we propose creating AI-Agents with personalities through a psychometric approach by leveraging Big Five Personality theory. Historically, Big Five personality theory was developed based on the Lexical Hypothesis, which states that personality characteristics fundamental to humans have become a part of human language, and the most important characteristics are encoded by a single word [caprara2000personality]. Based on the Lexical Hypothesis, scholars selected personality-related words from the English dictionary [allport1936trait], refined the list of words [norman1963toward, norman19672800, wiggins1979psychological], had participants rate themselves against those descriptors, applied factor analysis to reduce the dimensionality of the data, and eventually identified the Big Five Personality traits [mccrae1994openness, goldberg2013alternative, fiske1949consistency, peabody1989some]: Openness (O), Conscientiousness (C), Extraversion (E), Agreeableness (A) and Neuroticism (N).
Over the past decades, many psychometric scales [^1] The terms “scale” and “test” are used interchangeably to refer to psychometric instruments for assessing personality traits. measuring the Big Five traits have been developed, notably the Big Five Inventory [[, BFI;]]john_big_1991, Big Five Inventory-2 [[, BFI2;]]soto_next_2017, Mini-Markers [saucier_mini-markers_1994], the Big Five Aspects Scale [[, BFAS;]]deyoung_between_2007 and the NEO Five-Factor Inventory [[, NEO-FFI;]]costa_neo_1989. Despite having different items, these psychological measurements demonstrated high reliability, convergent validity and predictive validity in samples from diverse backgrounds [mccrae1997personality, mccrae2009physics, poortinga2002cross]. Specifically, these psychological measurements of the Big Five can significantly predict life outcomes, such as educational achievement, socioeconomic status, health, interpersonal relationships with peers and family, and more, indicating that personality traits can indeed forecast a range of individual, interpersonal and societal behaviors effectively [soto_how_2019, ozer2006personality, roberts2007power].
@@
Bringing psychometric principles to design and evaluate has practical benefits [wang_evaluating_2023, ye2025largelanguagemodelpsychometrics]. Integrating well-established psychometric tests (i.e., high reliability and validity) into the design of AI-Agents could ensure that the personality traits assigned to AI-Agents are stable and reflective of their programmed characteristics, enhancing the credibility and robustness of findings in applications where the consistency of AI-Agents performance is critical. In addition, as psychometric tests are designed to discern different levels of expression related to psychological traits, we can reverse engineer that variability into AI-Agents to create diversity, which is necessary for simulating real-world settings. Additionally, free from external confounds, this approach allows for precise and fine-grained manipulation over the psychological constructs of interest, allowing nuanced understanding and refinement of AI-Agents’ personalities for tailored applications.
@@ 2.1 Methods
Our analysis incorporated content from widely recognized Big Five tests, such as the BFI [john_big_1991], BFI-2 [soto_next_2017], Mini-Markers [saucier_mini-markers_1994], BFAS [deyoung_between_2007] and NEO-FFI [costa_neo_1989]. We extracted domain-specific content (i.e., test items) from these tests and processed it through OpenAI’s advanced text-embedding model (text-embedding-3-large). Text embeddings are commonly used in natural language processing; they allow researchers to represent text data (such as words, phrases, sentences, or even entire documents) in a numeric format (i.e., vector) that quantifies the semantic relationships between words, phrases, or entire documents. In the context of our study, each personality test item is converted into a high-dimensional vector (3072 dimensions for the model we used) that captures its semantic meaning.
@@ 2.2 Embeddings Techniques
We applied t-SNE [van2008visualizing] to item embeddings to create a two-dimensional visualization of their similarity structure. t-SNE is a non-linear dimensionality reduction method that maps high-dimensional data to a low-dimensional space while preserving local relationships between data points. In our context, the t-SNE provides an intuitive visual map of semantic neighbourhoods among personality test items, which provides a good complement to the cosine similarity analysis. Items from the same Big Five domain tend to cluster together in the plot, and cross-domain proximities become easier to spot visually. In our t-SNE visualizations, items that cluster together can be interpreted as measuring similar constructs [^2] Because t-SNE is intended for visualization rather than formal measurement modelling, the plots are used here descriptively to illustrate patterns in the embedding space; the axes have no intrinsic psychological meaning..
@@ 3.1.1 Method
We included four prompt conditions for creating AI-Agents with personalities (see Table 1). Two conditions involved using the BFI-2 in either Likert or Expanded format, which we refer to as BFI-2-Likert and BFI-2-Expanded, respectively. The BFI-2-Likert is the original BFI-2 by [soto_next_2017], in which participants indicated their agreement with statements measuring the Big Five traits. The BFI-2-Likert, however, is susceptible to response bias and careless responding; therefore, [Zhang_Huang_2025] developed the BFI-2-Expanded version, replacing the agree-disagree response options with complete sentences. For example, for the Expanded format, the “strongly agree” option for the item “Is outgoing, sociable” is replaced with “I am very outgoing, sociable”. For prompt engineering, the Expanded format produces prompts that sound more natural and straightforward. As shown in Table 1, the BFI-2-Likert prompt requires an additional instruction for interpreting numeric numbers indicating agreement levels, whereas the BFI-2-Expanded prompt simply describes the trait levels in complete sentences.
We also included two baseline prompt conditions (Simple-Binary and Elaborated-Binary), designed to mimic the approach of previous studies that used simple adjectives to create AI-Agent personas [jiang_personallm_2024, park_generative_2023, serapio2023personality]. In the Simple-Binary condition, prompts were binary Big Five statements indicating whether the AI-Agent was high or low on each trait. The Elaborated-Binary condition added descriptors such as “outgoing” and “compassionate” to further elaborate on the traits. Compared with these baseline conditions, we aim to show that the BFI-2-Likert or BFI-2-Expanded condition create AI-Agents with more nuanced personalities and aligns better with human responses.
@@
### Your Personality ### I am very outgoing, sociable. I am very compassionate, almost always soft-hearted. I am fairly organized. I am somewhat relaxed, handle stress somewhat well. I have some artistic interests. I am quite assertive. I am very respectful almost always treat others with respect. I am often lazy. I stay very optimistic after experiencing a setback. I am curious about few things. I often feel excited or eager. I rarely find fault with others. I am very dependable, steady. I am fairly moody often have up and down mood swings. I am fairly inventive, often find clever ways to do things. I am rarely quiet. I feel a great deal of sympathy for others. I am not particularly systematic rarely keep things in order. I am sometimes tense. I am very much fascinated by art, music or literature. I am fairly submissive, often act as a follower. I rarely start arguments with others. I have a fair amount of difficulty getting started on tasks. I feel very secure, comfortable with self. I typically seek out intellectual, philosophical discussions. I am somewhat more active than other people. I have a very forgiving nature. I am often careless. I am fairly emotionally stable quite hard to upset. I have little creativity. I am often shy, introverted. I am very helpful and unselfish with others. I sometimes keep things neat and tidy. I worry quite a lot. I value art and beauty quite a bit. I find it fairly easy to influence people. I am sometimes rude to others. I am fairly efficient, get things done fairly quickly. I almost never feel sad. I am not particularly complex rarely a deep thinker. I am almost always full of energy. I am quite trusting of others’ intentions. I am fairly reliable, can usually be counted on. I usually keep my emotions under control. I have a bit of difficulty imagining things. I am very talkative. I am very warm and caring. I rarely leave a mess usually clean up. I often feel anxious or afraid. I think poetry and plays are fairly interesting. I strongly prefer to take charge. I am very polite and courteous to others. I am fairly persistent, usually work until the task is finished. I sometimes feel depressed and blue. I have little interest in abstract ideas. I show a lot of enthusiasm. I almost always assume the best about people. I often behave irresponsibly. I am not at all temperamental almost never get emotional. I am not at all original almost never come up with new ideas.
@@
To examine alignment between AI-Agents responses with those from human participants, we repurposed data collected by [soto_next_2017] wherein participants (N = 438) responded to multiple Big Five tests, including: (1) BFI-2: a sophisticated 60-item Likert test designed to capture the comprehensive hierarchical structure of personality, where each domain has three facets and (2) Mini-Markers: a straightforward, 40-item test consisting of phenotypic Big Five descriptive adjectives [saucier_mini-markers_1994]. Participants’ BFI2 responses will serve as the training (input) data, while their Mini-Markers responses (output) will be used for validation. The human data [soto_next_2017] will serve as a reference for our comparison.
@@
Using the four types of prompts in Table 1, we created AI-Agents across five popular LLMs, including 1) GPT-3.5-Turbo (01-25), 2) GPT-4-Turbo (04-09), 3) GPT-4o (2024-11-20), 4) DeepSeek-V3 (2024-12-26), and 5) Llama-3.3-70B-Instruct (2024-12-06). For each LLM and each prompt condition, we initiated 438 AI-Agents, each corresponding to a human participant’s data collected from [soto_next_2017]. More specifically, for the BFI-2-Likert and BFI-2-Expanded conditions, AI-Agents were created by inputting prompts that matched the BFI-2 scores of the corresponding human participants. For the Baseline-Simple and Baseline-Elaborated conditions, participants’ trait scores were dichotomized as high (average score >2.5>2.5) or low (average score <2.5<2.5) for each of the Big Five traits, and AI-Agents were then created using these dichotomized scores. Across all conditions, the LLM temperature was set to the default value of 1.0 to balance coherence and diversity, and other sampling parameters (e.g., top-p, frequency, and presence penalties) were kept at their defaults.
After creating AI-Agents for each LLM and prompt condition, we prompted them to complete the Mini-Markers test, which are 40 adjectives describing the Big Five traits [saucier_mini-markers_1994]. The full prompt for this task is provided in the Appendix A. In this prompt, AI-Agents were instructed to consider the Mini-Marker items carefully based on their assigned personality (see Appendix).